# DeFiScope: Detecting Various DeFi Price Manipulations with LLM Reasoning

### v1

@@ 7.2. RQ2: Ablation Study
In this RQ, we investigate how fine-tuning can enhance DeFiScope’s detection accuracy on the same D1 dataset, as well as the impact and cost of fine-tuning different GPT models. To this end, we test four settings shown in Figure 9: (a) the original GPT-3.5-Turbo without fine-tuning, (b) GPT-3.5-Turbo with fine-tuning, which was used in RQ1, (c) the original GPT-4o without fine-tuning, and (d) GPT-4o with fine-tuning.
@@
Figure 9 shows that fine-tuning significantly enhances DeFiScope’s ability to detect attacks, with the fine-tuned versions of GPT-3.5-Turbo and GPT-4o detecting 18 (31%) and 12 (19%) more attacks, respectively. It also indicates that fine-tuning provides a more noticeable improvement for less powerful models, such as GPT-3.5-Turbo, compared to stronger models like GPT-4o. With fine-tuning, the detection success rate for attacks targeting the CPMM increases to 100% with GPT-3.5-Turbo and 95.6% with GPT-4o [^4] GPT-4o does not exhibit an advantage with CPMM but shows a clear advantage with custom price models, especially when the original GPT-4o compared to GPT-3.5-Turbo.. This very high rate can be attributed to the use of CPMM during the fine-tuning training phase, enabling the model to effectively handle this specific pricing model.
@@
Figure 9. Effectiveness of fine-tuning for 78 attacks that go through DeFiScope’s LLM inference. The cases not plotted in the figures include eight due to missing source code and compilation errors, and nine due to cross-transaction issues and non-adherence to the ERC20 token standard, as mentioned in §7.1.
@@
Table 5. Comparison between GPT-3.5-Turbo and GPT-4o.
Model $TP$ $Recall$ Ave. Cost(\$)/ per inference Fine-tuning Cost(\$)
@@
LLM Costs. Table 5 highlights that while the original GPT-4o generally performs better than GPT-3.5-Turbo, this performance gap can be largely minimized through fine-tuning. However, regarding costs, GPT-3.5-Turbo has a clear advantage, which is why DeFiScope uses GPT-3.5-Turbo for fine-tuning by default instead of GPT-4o. Specifically, the fine-tuning cost of GPT-3.5-Turbo is 68% lower than that of GPT-4o, and the per-request inference cost of GPT-3.5-Turbo is merely \$0.0107 (OpenAi, 2024b), which is also 18% cheaper than that of GPT-4o. Therefore, for a tradeoff between cost and performance, we recommend using GPT-3.5-Turbo for DeFiScope.
7.3. RQ3: Real-World Practicality
@@
To be practical in a real-world setting, DeFiScope not only needs to maintain high detection rates for true attacks but also minimize false alarms for benign transactions. In this RQ, we evaluate this aspect of DeFiScope and the associated time overhead using the datasets D2 and D3, which were introduced in the prologue of §7. Specifically, D2 comprises 968 suspicious transactions with various attacks (i.e., not limited to price manipulations), and D3 includes 96,800 benign transactions collected from DeFort (Xie et al., 2024)’s dataset.
@@
For D2, DeFiScope flagged 153 price manipulation attacks out of 968 suspicious transactions. To robustly confirm these potential attacks, we cross-referenced them with attack reports or alerts published by security companies through their official channels (Ancilia, 2024a; Beosin, 2024; BlockSec, 2024b; SlowMist, 2024b), thereby verifying the root causes of the attacks. Using this method, we confirm that 66 of them are previously reported price manipulation attacks. For the remaining cases, we conducted comprehensive and in-depth analyses by combining manual review with ancillary evidence, such as identifying whether the EOA initiating the transaction was marked as a hacker by blockchain explorers (BscScan, 2024; EtherScan, 2024; OKLink, 2024). Finally, we discovered 81 previously unknown historical incidents, which were neither reported by security companies nor tagged as malicious transactions by blockchain explorers, and identified a total of six false positive cases (as detailed in Table 6 in Appendix C). Of these, five are benign transactions initiated by the same EOA and exhibiting identical invocation flows, while the last one is indeed an attack due to a logic issue rather than a price manipulation attack, resulting in a precision of 96% (147/153).
In comparison, DeFort identified only 58 attacks, of which 9 were false positives labeled as other vulnerabilities by security researchers. Moreover, DeFort failed to detect 114 of the malicious transactions flagged by DeFiScope, including 54 that are officially confirmed attacks.
@@
We further analyze DeFiScope’s false alarms in a realistic setting mixing 968 suspicious transactions in D2 with 96,800 benign transactions in D3 (see the earlier dataset setup in the prologue of this section). For each transaction, we set the maximum scan time to 300 seconds. The results reveal that DeFiScope achieves zero false alarms on benign transactions in D3, with an average of 2.5 seconds per transaction across all transactions. This highlights DeFiScope’s potential for large-scale, daily on-chain monitoring scenarios.
8. Discussion
While the precision of DeFiScope is quite satisfactory, its recall rate primarily affected by the following two threats. Future work can further optimize the following two aspects:
Cross-transaction price manipulation attacks. DeFiScope currently detects only single-transaction attacks. However, some price manipulation attacks are designed across multiple transactions to circumvent time restrictions coded in DeFi protocols, as studied in (Chen et al., 2024b). Detecting cross-transaction attacks is often challenging. For example, attacks like INUKO (Ancilia, 2024b) span 48 hours (crossing around 57,000 blocks), making it extremely difficult to precisely identify all related transactions. Fortunately, our study observed that cross-transaction attacks are less common than single-transaction ones.
@@ 8. Discussion
Closed-source price calculation functions. The availability of code for price models could affect the detection accuracy of DeFiScope. According to our study, most DeFi applications are open-source to gain user trust. To mitigate issues with closed-source price models, we design the Type-II prompt to cover price models in those closed-source liquidity pools.
9. Related Work
@@
On-chain Security Analysis. Prior efforts (Zhou et al., 2021b; Qin et al., 2021, 2022; Daian et al., 2020; Zhou et al., 2023) have elucidated threats in blockchain and attacks against the DeFi ecosystem. Some transaction-based systems have been proposed to mine vulnerable transaction sequences (Zhang et al., 2020; So et al., 2021), explore arbitrage opportunities (Zhou et al., 2021a), detect malicious phishing (He et al., 2023), and simulate attacks to prevent intrusions (Qin et al., 2023a).
Tools like FlashSyn (Chen et al., 2024a), DeFiRanger (Wu et al., 2023), DeFiTainter (Kong et al., 2023), and DeFort (Xie et al., 2024) are capable of detecting attacks associated with price manipulation. FlashSyn uses numerical approximation techniques to synthesize malicious contracts that target DeFi apps through price manipulation attacks. However, there is a gap between the synthesized contracts and those that launch attacks on the live blockchain.DeFiRanger generates cash flow trees (CFT) from a transaction, translates the semantics of CFT into high-level DeFi actions, and uses specific patterns to detect price manipulation attacks. Yet, DeFiRanger is significantly limited by its reliance on predefined templates. DeFiTainter constructs call graphs based on contract states and transactions, then performs inter-contract taint analysis to examine all execution paths and uncover price manipulation attacks. Nonetheless, DeFiTainter’s approach requires manual analysis and labeling of taint sources and sinks, which is challenging. DeFort uses a price manipulation behavior model to guide detection and analysis but its boundary calculation mechanism lacks sensitivity to attacks characterized by minor price fluctuations.
Program Analysis for Contract Security. Traditional tools such as Slither (Feist et al., 2019), Zeus (Kalra et al., 2018), and Securify (Tsankov et al., 2018) employ static analysis methods to detect vulnerabilities at the source code and bytecode levels. Symbolic execution has also shown significant potential in this area, with tools like Mythril (Consensys, 2024), Oyente (Luu et al., 2016), and Manticore (Mossberg et al., 2019) developed based on this technique. Dynamic analysis is another core methodology utilized by various tools (Jiang et al., 2018; Liu et al., 2018; Shou et al., 2023; Qin et al., 2023b). Some tools are specifically designed to detect particular types of vulnerabilities; for example, VetSC (Duan et al., 2022) identifies business logic bugs by vetting inconsistencies between UI and contract code, SailFish (Bose et al., 2022) focuses on state inconsistency issues, and JACKAL (Gritti et al., 2023) targets confused deputy vulnerabilities.
@@
LLMs for Smart Contract Security. LLMs have become formidable tools in the realm of blockchain security. GPTScan (Sun et al., 2024b) uses LLMs for code understanding, incorporating static analysis to detect logical vulnerabilities in contracts. BlockGPT (Gai et al., 2023) serves as a real-time intrusion detection system, using LLMs to rank transaction anomalies. LLM4Vuln (Sun et al., 2024a) refines the reasoning processes of LLMs when analyzing smart contract vulnerabilities. iAudit (Ma et al., 2025) combines fine-tuning with LLM-driven agents to provide intuitive smart contract auditing with explanations. PropertyGPT (Liu et al., 2025) uses retrieval-augmented property generation for smart contracts, leveraging LLMs’ in-context learning capabilities. Similarly, SMARTINV (Wang et al., 2024) employs Tier of Thought (ToT) to reason across different modalities of smart contracts, generating invariants in the process. To the best of our knowledge, DeFiScope is the first tool that uses LLMs specifically designed for detecting price manipulation attacks.
10. Conclusion
@@
Appendix
Table 6. Details of six false positives mentioned in §7.3.
Transaction hash Chain Block Type Root Cause
0x130c6370 BSC 38218540 Benign -
0x4b59af93 BSC 38218538 Benign -
0xe158a2b9 BSC 38218537 Benign -
0x59942848 BSC 38218536 Benign -
0x2e9ceb16 BSC 38218539 Benign -
0x640ce34c BSC 11403670 Malicious Logic issue
Appendix A A Case Study of Type-II Prompt
Figure 10 illustrates an use case of Type-II prompt. The contract (0x2120…3379) is closed source DEX contract that allows users to trade SVT and USDT. While recovering the DeFi operations, this contract was identified as a two-token liquidity pool. So, we applied the Type-II prompt to inferring the token price trend in its transactions. As shown by the LLM response in Figure 10, the primary difference from the response to a Type-I prompt lies in the absence of the analysis on the price model; instead, the scoring is directly yielded. The is because the prompt itself assumes that the CPMM is employed in this contract. Regarding the CPMM, the relationship between token price and token balance is quite standardized, i.e., the direction of balance change is opposite to that of price change.
Figure 10. An example of Type-II prompt and its response generated while detecting the attack against SVT (BlockSec, 2024a).
Table 7. High-value DeFi applications (from DefiLlama (DefiLlama, 2024)) with active transactions of each category we have done a comprehensive study on. TVL refers to the Total Value Locked in the app.
DEX App TVL Lending App TVL Yield farming App TVL
Uniswap \$4.8B Compound \$2.0B Pendle \$2.8B
Curve \$1.9B AAVE \$12.3B Convex Finance \$1.1B
Balancer \$777.2M Morpho \$1.5B Aura \$361.9M
Sushi \$250.2M Fraxlend \$134.4M Magpie \$192.5M
PancakeSwap \$1.7B Venus \$1.5B StakeDAO \$78.8M
1inch \$4.58M Strike \$9.7M Equilibria Finance \$80.0M
ParaSwap \$6.34M Planet \$1.2M Kine Finance \$8.0M
ShibaSwap \$18.19M Kinza Finance \$42.8M Dot Dot Finance \$2.1M
BiSwap \$27.1M Radiant \$7.8M Solo Top \$1.8M
MDEX \$16.0M Ambit \$6M Jetfuel Finance \$1.5M
Appendix B High-value DeFi Applications
The design of DeFi operations and the category of transfer actions in §5 is based on an in-depth study of the high-value DeFi applications listed in Table 7 as of August 2024.
Appendix C Details of False Positives
Table 6 presents the details of six false positives discovered in our experiments described in §7.3. The reason for these false positives is that their transactions involve two contract accounts that were created three months ago by the transaction initiator. Any fund transfers among these accounts and the initiator should be considered benign operations rather than price manipulation operations. Yet, these accounts were incorrectly marked as closed-source DEXes in the detection, leading to false inferences. Such false positives could be mitigated by conducting a historical analysis of account ownership relationships and clustering user accounts that are controlled by the same owners.

### v2

@@ VII-B RQ2: Ablation Study
In this RQ, we investigate how fine-tuning can enhance DeFiScope’s detection accuracy on the same D1 dataset, as well as the impact and cost of fine-tuning different GPT models. To this end, we test four settings shown in Fig. 6: (a) the original GPT-3.5-Turbo without fine-tuning, (b) GPT-3.5-Turbo with fine-tuning, used in RQ1, (c) the original GPT-4o without fine-tuning, and (d) GPT-4o with fine-tuning.
@@
Fig. 6 shows that fine-tuning significantly enhances DeFiScope’s ability to detect attacks, with the fine-tuned versions of GPT-3.5-Turbo and GPT-4o detecting 18 (31%) and 12 (19%) more attacks, respectively. It also indicates that fine-tuning provides a more noticeable improvement for less powerful models, such as GPT-3.5-Turbo, compared to stronger models like GPT-4o. With fine-tuning, the detection success rate for attacks targeting CPMM increases to 100% with GPT-3.5-Turbo and 95.6% with GPT-4o [^2] GPT-4o does not exhibit an advantage with CPMM but shows a clear advantage with custom price models, especially for the original GPT-4o.. This very high rate can be attributed to the use of CPMM during the fine-tuning training phase, enabling the model to effectively handle it.
@@
Figure 6: Effectiveness of fine-tuning for 78 attacks that go through DeFiScope’s LLM inference. Cases not plotted in the figures include eight due to missing source code and compilation errors, and nine due to cross-transaction issues and non-adherence to the ERC20 token standard, as mentioned in §VII-A.
@@
LLM Costs. Table V highlights that while the original GPT-4o generally performs better than GPT-3.5-Turbo, this performance gap can be largely minimized through fine-tuning. However, regarding costs, GPT-3.5-Turbo has a clear advantage, which is why DeFiScope uses GPT-3.5-Turbo for fine-tuning by default instead of GPT-4o. Specifically, the fine-tuning cost of GPT-3.5-Turbo is 68% lower than that of GPT-4o, and the per-request inference cost of GPT-3.5-Turbo is merely \$0.0107 [60], which is also 18% cheaper than that of GPT-4o. Therefore, for a tradeoff between cost and performance, we recommend using GPT-3.5-Turbo.
TABLE V: Comparison between GPT-3.5-Turbo and GPT-4o.
Model TPTP RecallRecall Ave. Cost(\$)/ per inference Fine-tuning Cost(\$)
@@
VII-C RQ3: Real-World Practicality
@@
To be practical in a real-world setting, DeFiScope not only needs to maintain high detection rates for true attacks but also minimize false alarms for benign transactions. In this RQ, we evaluate this aspect of DeFiScope and the associated time overhead using the datasets D2 and D3, which were introduced in the prologue of §VII. Specifically, D2 comprises 968 suspicious transactions with various attacks (i.e., not limited to price manipulations), and D3 includes 96,800 benign transactions collected from DeFort [5]’s dataset.
@@
For D2, DeFiScope flagged 153 price manipulation attacks out of 968 suspicious transactions. To robustly confirm these potential attacks, we cross-referenced them with attack reports or alerts published by security companies through their official channels [61, 62, 63, 64], thereby verifying the root causes of the attacks. Using this method, we confirm that 66 of them are previously reported price manipulation attacks. For the remaining cases, we conducted comprehensive and in-depth analysis by combining manual review with ancillary evidence, such as identifying whether the EOA initiating the transaction was marked as a hacker by blockchain explorers [65, 66, 67]. For newly discovered incidents, we manually verified: (i) whether abnormal token price or exchange rate changes occured, and (ii) whether the user gained profit from such changes. Cases satisfying both criteria were labeled as price manipulation attacks. Finally, we discovered 81 previously unknown historical incidents, which were neither reported by security companies nor tagged as malicious transactions by blockchain explorers, and identified a total of six false positive cases (the full list can be found in [28]). Of these, five are benign transactions initiated by the same EOA and exhibiting identical invocation flows, while the last one is indeed an attack due to a logic issue rather than a price manipulation attack, resulting in a precision of 96% (147/153) on suspicious transactions. In comparison, DeFort identified only 58 attacks, of which 9 were false positives labeled as other vulnerabilities by security researchers. Moreover, DeFort failed to detect 114 of the malicious transactions flagged by DeFiScope, including 54 that are officially confirmed attacks.
@@
We further analyze DeFiScope’s false alarms in a realistic setting mixing 968 suspicious transactions in D2 with 96,800 benign transactions in D3 (see the earlier dataset setup in the prologue of this section). For each transaction, we set the maximum scan time to 300 seconds. The results reveal that DeFiScope achieves zero false alarms on benign transactions in D3, and 6 false positives on D2 (as mentioned above), with an average of 2.5 seconds per transaction across all transactions. More than half the time is spent on price change inference that uses up an average of 1.40 seconds, while static analysis takes slightly less, around 0.97 seconds on average. The other steps within DeFiScope incur negligible cost. A detailed time cost breakdown is available in  [28]. This highlights DeFiScope’s potential for large-scale, daily on-chain monitoring scenarios.
VIII Discussion
Cross-LLM Generalizability. To show DeFiScope’s methodology is not limited to OpenAI’s fine-tuning paradigm [22], we performed SFT fine-tuning [23] on a open-source Phi-3 [68] model using LoRA [24], on a sever with 4 Nvidia H800 GPUs. We used all 1,000 price change samples mentioned in §IV, split 8:1:1 into training, evaluation, and test sets. As shown in Table VI, LoRA-based fine-tuning enabled Phi-3 to outperform GPT-3.5-Turbo (without fine-tuning) and achieve performance comparable to GPT-4o (without fine-tuning). Full hyperparameter settings and results are available in our repository [69].
TABLE VI: Comparison between GPT tuning and Phi-3’s SFT.
Model RecallRecall Model RecallRecall
GPT-3.5 w/o fine-tuning 0.61 GPT-3.5 w/ fine-tuning 0.80
GPT-4o w/o fine-tuning 0.66 GPT-4o w/ fine-tuning 0.79
Phi-3 w/o fine-tuning 0.52 Phi-3 w/ fine-tuning [70] 0.66
Cross-Blockchain Generalizability. Although our evaluation primarily focused on Ethereum and BSC, which together account for the majority of observed price manipulation attacks, DeFiScope is generalizable to EVM-compatible blockchains, such as Polygon [71] and Arbitrum [72]. DeFiScope successfully detected seven price manipulation attacks on these chains. Full details of these cases are available in [28].
@@ VIII Discussion
Closed-source contracts and cross transaction attacks. The availability of price model code can affect DeFiScope’s detection accuracy. According to our study, most DeFi applications are open-source to gain user trust. To mitigate issues with closed-source price models, we design the Type-II prompt to cover price models in those closed-source liquidity pools. Currently, DeFiScope detects only single-transaction attacks, while some attacks span multiple transactions to bypass protocol time restrictions [73]. Detecting such cases is challenging. For instance, attacks like [48] span over 48 hours (about 57,000 blocks), making it difficult to trace all related transactions. Fortunately, our study observed that such attacks are far less common than single-transaction ones.
IX Related Work
@@
On-chain Security Analysis. Prior efforts [74, 75, 76, 77, 78] have studied blockchain threats and DeFi attacks. Some transaction-based systems are proposed to mine vulnerable transaction sequences [79, 80], explore arbitrage opportunities [81], detect malicious phishing [82], and simulate attacks to prevent intrusions [83]. Tools like FlashSyn [45], DeFiRanger [4], DeFiTainter [7], and DeFort [5] are capable of detecting attacks associated with price manipulation. FlashSyn uses numerical approximation to synthesize malicious contracts that target DeFi apps through price manipulation attacks.
@@
LLMs for Smart Contract Security. LLMs have become powerful tools in blockchain security. GPTScan [84] leverages LLMs with static analysis to detect logical vulnerabilities. BlockGPT [85] uses LLMs to rank transaction anomalies for real-time intrusion detection. LLM4Vuln [86] improves LLM reasoning for vulnerability analysis. iAudit [25] combines fine-tuning with agents for intuitive auditing. PropertyGPT [87] uses LLMs for retrieval-augmented property generation. To our knowledge, DeFiScope is the first tool that uses LLMs specifically designed for detecting price manipulation attacks.
X Conclusion
@@
[…]
Acknowledgment We thank all reviewers for their constructive feedback. This research was supported by Lingnan Grant SUG-002/2526, HKUST TLIP Grant FF612, the National Natural Science Foundation of China (Project No. 72304232), the Singapore Ministry of Education Academic Research Fund Tier 2 (T2EP20224-0003) and the Nanyang Technological University Centre for Computational Technologies in Finance (NTU-CCTF). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of MOE and NTU-CCTF.